An electrical firm manufactures light bulbs that have a length of life that is normally distributed with mean equal to 778 hours and a standard deviation of 40 hours. Find the probability that a bulb burns between 710 and 841 hours.


1
Expert's answer
2022-02-07T17:44:13-0500

Let "X=" a length of life: "X\sim N(\mu, \sigma^2)."

Given "\mu=778\ h, \sigma=40\ h."


"P(710<X<841)=P(X<841)-P(X\le 710)"

"=P(Z<\dfrac{841-778}{40})-P(Z\le \dfrac{710-778}{40})"

"= P(Z<1.575)-P(X\le -1.7)"

"\approx0.94237-0.04457= 0.8978"